# Stability of Non-Hyperbolic Sets

### v1

@@ 12 Properties of the Lamination: Collet-Eckmann and Newhouse points
The following Theorem proves Theorem B.
Theorem 7.
There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$. Moreover the closure $\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
@@
The proof goes by induction on the generations $g$ defining the lamination. The induction step is contained in Proposition 8. Using the notation in that proposition we prove the base of induction, i.e., we prove that the conditions are verified at generation $g=1$.
Consider the $1$-tangency strip $S$ given by $S=[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with tangency line $b=\left\{a=0\right\}$. Take $U=S$. In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2, property $(f10)$. Define $T=0$ and for all $f\in U$ let $K$ be a positive constant assuring that $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(iii) and $\text{IH}_{C^{N}}$-(iii) hold. The existence of such a constant is guaranteed by the initial conditions on the family. Observe that $\text{IH}_{Q}$, $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(i) and $\text{IH}_{C^{0}}$-(ii) are consequence of the definition of our family, while $\text{IH}_{C^{N}}$ is guaranteed since $T=0$. As for $\text{IH}_{C^{0}}$-(ii), the lower bound comes from the fact that $b$ is a non-degenerate tangency, while the upper bound is a consequence of the fact that the family is made of ${\mathcal{C}}^{2}$} diffeomorphisms. This concludes the base of induction. As said above, the induction step is Proposition 8. We can now draw our conclusions. Let
@@
Since $T_{g}$ goes to infinity, $c$ is a Collet-Eckmann point. The proof of the density is the same as the proof of Lemma 24. ∎
Even more surprisingly, there exists a sub-lamination of $\mathfrak{L}_{CE}$ whose leaves contains a map with infinitely many sinks of higher and higher order accumulating at the orbit of the Collet-Eckmann point. The following Theorem proves Theorem C.
Theorem 8.
There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. Each leave $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{k}$ of arbitrarily high period, such that,
$\bigcap_{n}\overline{\bigcup_{k\geq n}\text{Orb}(p_{k})}=A(f_{\ell}).$
Moreover if $NC=\left\{f_{\ell}\right\}_{f_{\in}\ell\subset\mathfrak{L}_{N}}$, then the closure $\overline{{NC}}=\overline{\mathfrak{L}_{CE}}=\overline{\mathfrak{L}}$.
@@
The theorem is consequence of Theorem 7 and Theorem 3. ∎
@@ 13 Appendix
We prove that there exists a $\theta$ satisfying (8.1).
Lemma 30.
@@
By (4.17), there exists $K>1$ such that
$\alpha=\frac{\log K}{\log\mu}+2\theta\frac{\log\lambda}{\log\mu}+3.$
@@
$\log K\left(1-\frac{\log\lambda}{\log\mu}\right)-3\log\lambda+4\log\mu>\theta\left(2\frac{(\log\lambda)^{2}}{\log\mu}-3\log\lambda\right).$

### v3

@@ 12 Collet-Eckmann Lamination and Newhouse points
The following Theorem confirms Theorem B.
Theorem 12.0.2.
There exists a sub-lamination $\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. For each function $f\in\ell\subset\mathfrak{L}_{CE}$, the invariant non-hyperbolic Cantor set $A(f)$ contains a Collet-Eckmann point $c$ such that $\omega(c)=A(f)$. Moreover, every leaf of $\mathfrak{L}_{CE}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{CE}_{t}=\left\{(t,a)\in\mathfrak{L}_{CE}\right\}$ is a Cantor set.
@@
The proof goes by induction on the generations $g$ defining the lamination. The induction step is contained in Proposition 8.0.4. Using the notation in that proposition we prove the base of induction, i.e., we prove that the conditions are verified at generation $g=1$.
Consider the $1$-tangency strip $S$ given by $S=[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with tangency line $b=\left\{a=0\right\}$. Take $U=S$. In Subsection 2.1, we defined a family $F$ which is an unfolding of the tangency $b$ and each map $f\in S$ has a $(1,N_{1},M_{1})$ dynamical partition with $N_{1}=N$ and $M_{1}=M$, where $N$ and $M$ are defined in Definition 2.1.1, property $(f8)$. Define $T=0$ and for all $f\in U$ let $K$ be a positive constant assuring that $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(iii) and $\text{IH}_{C^{N}}$-(iii) hold. The existence of such a constant is guaranteed by the initial conditions on the family. Observe that $\text{IH}_{Q}$, $\text{IH}_{H^{0}}$, $\text{IH}_{C^{0}}$-(i) and $\text{IH}_{C^{0}}$-(ii) are consequence of the definition of our family, while $\text{IH}_{C^{N}}$ is guaranteed since $T=0$. As for $\text{IH}_{C^{0}}$-(ii), the lower bound comes from the fact that $b$ is a non-degenerate tangency, while the upper bound is a consequence of the fact that the family is made of ${\mathcal{C}}^{2}$} diffeomorphisms. This concludes the base of induction. As said above, the induction step is Proposition 8.0.4. We can now draw our conclusions. Let
@@
Since $T_{g}$ goes to infinity, $c$ is a Collet-Eckmann point. The proof of the density is the same as the proof of Lemma 9.3.4. ∎
Even more surprisingly, there exists a sub-lamination of $\mathfrak{L}_{CE}$ whose leaves contains a map with infinitely many sinks of higher and higher order accumulating at the orbit of the Collet-Eckmann point. The following Theorem confirms Theorem C.
Theorem 12.0.3.
There exists a sub-lamination $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}\subset\mathfrak{L}$ satisfying the following property. Each leaf $\ell\subset\mathfrak{L}_{N}$ contains a map $f_{\ell}$ with infinitely many attracting periodic points, $p_{g}$ of arbitrarily high period, such that,
$\bigcap_{n}\overline{\bigcup_{g\geq n}\text{Orb}(p_{g})}=A(f_{\ell}).$
Moreover, every leaf of $\mathfrak{L}_{N}$ is the graph of a smooth function $\ell:[-t_{0},t_{0}]\mapsto[-a_{0},a_{0}]$ and if $t\in[-t_{0},t_{0}]$, the closure of the set $L^{N}_{t}=\left\{(t,a)\in\mathfrak{L}_{N}\right\}$ is a Cantor set.
@@
The theorem is a consequence of Theorem 12.0.2 and Theorem 6.2.7. More in details, using the terminology from Subsection 6.2, given a point $(t,a)\in NH$ there exists a sequence of Newhouse boxes such that
$\left\{(t,a)\right\}=\bigcap_{g\geq 1}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
Associated to each Newhouse box $\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}$ there is a secondary tangency curve which is the graph of $b^{g}_{n^{(g)},n_{0}^{(g)}}$. By taking $n^{(g)}$ large enough we can assure that these curves converge in ${\mathcal{C}}^{1}$} to a leaf $\ell_{(t,a)}$ which is also the graph of a ${\mathcal{C}}^{1}$} function over $t_{0},t_{0}]$. In particular $(t,a)\in\ell_{(t,a)}$. Let
$\mathfrak{L}_{N}=\bigcup_{(t,a)\in NH}\ell_{(t,a)}.$
By taking $n^{(g)}$ large enough we can assure that $\mathfrak{L}_{N}\subset\mathfrak{L}_{CE}$.
Let $p_{g}$ be the periodic sink of generation $g$. By construction the orbit of $p_{g}$ is contained in the $c$- and $q$-loop of the dynamical partition $\mathfrak{D}_{g}$. Hence, the sinks accumulate at the Cantor set $A$. ∎
@@ 13 Appendix
We prove that there exists a $\theta$ satisfying (8.0.2).
Lemma 13.0.1.
@@
By (4.1.47),
$\alpha=2\theta\frac{\log\lambda}{\log\mu}+3.$
@@
$3\log\lambda+4\log\mu>\theta\left(2\frac{(\log\lambda)^{2}}{\log\mu}-3\log\lambda\right).$